OpenStudy (anonymous):

Can someone help me to solve the already intergrated problem so I can compare my answer absolute (6t - 3t^2) from x = 1 to x=0 + absolute (6t -3t^2) from x= 5 to x= 1

OpenStudy (anonymous):

don't you mean from x=0 to x=2 and negative integral from x=2 to x=5 ... because the formula crosses x-axis at x=2.

OpenStudy (anonymous):

we got 6t-3t^3... integrating reveals: 3t^2-t^3. from x=0-2 4 from x= 2-5 -54 total area between graph and x-axis : 58